OpenStudy (anonymous):

A customer has an offer for 5 portrait sheets for 9.99 with additional portrait sheets priced at 19.99 each. If the customer desires to purchase 10 portrait sheets, how much would she save if she instead purchases a 10 portrait sheet package for $99.99?

hero (hero):

Thought I answered this one already

OpenStudy (anonymous):

yea my mistake

OpenStudy (anonymous):

$9.99 + 5 x $19.99 = 10 sheets total compared to: $99.99= 10 sheets total 5 x 19.99= 99.95 99.95+9.99= 109.94 109.94-99.99=9.95 You save $9.95